OpenStudy (anonymous):

for which value of k will the vector u=(1;-2;k) in R^3 be a linear combination of the vectors v=(3;0;-2) and w=(2;-1;-5)

OpenStudy (amistre64):

3 2 1 0 -1 * (c1,c2) = -2 -2 -5 k

OpenStudy (anonymous):

pls use matrix representation

OpenStudy (amistre64):

augment the left with the right into rref

OpenStudy (anonymous):

Just set it as a matrix right and row reduce it

OpenStudy (amistre64):

3 2 1 0 -1 -2 -2 -5 k yes

OpenStudy (amistre64):

another method is to determine for what values of k the determinant is zero

OpenStudy (amistre64):

-(2+30+0) 3 2 1 3 2 0 -1 -2 0 -1 -2 -5 k -2 -5 +(-3k+8+0)

OpenStudy (amistre64):

-3k+8-32 = 0 -3k = 24 k = -24/3 = -8

OpenStudy (phi):

I would just rref on the 3 x 3 matrix to get 3 0 -2 0 -1 -11/3 0 0 24/3 + k then say the last row must be all zeros. this means k= -8

OpenStudy (amistre64):

3 2 1 0 -1 -2 -2 -5 k 1 2/3 1/3 0 1 2 0 -11/3 (3k+2)/3 1 2/3 1/3 0 1 2 0 1 -(3k+2)/11 1 2/3 1/3 0 1 2 0 0 -(3k+2)-22/11 as long as the last row is zeros, this thing is "dependant" -(3k+2)-22/11 = 0 -3k-2-22 = 0 -3k-24 = 0 ... back to -8 again

OpenStudy (amistre64):

i personally think the determinant method is simpler

OpenStudy (anonymous):

Yeah but don't you learn that later....

OpenStudy (amistre64):

phi, yours is not rref, effect but not rref :) just an echelon form

OpenStudy (amistre64):

learn what later?

OpenStudy (phi):

@amistre64 i personally think the determinant method is simpler I think computationally its cost grows N! so for large N, you would want to do elimination. but for this problem, definitely!

OpenStudy (phi):

yes, just ref (not reduced row), sorry if that causes any confusion.

OpenStudy (anonymous):

determinants

OpenStudy (amistre64):

phi, right, i was blinded by the 3x3 of it ... lol

OpenStudy (amistre64):

Ryan, depends onwhat level Remainder is at i spose

OpenStudy (anonymous):

 can u pls check this i 've didn't learn determinat method

OpenStudy (anonymous):

@phi and @RyanL. CHECK THIS PLS!

OpenStudy (phi):

why did you make the 3rd column all zeros? if you are solving for Mx=b augment M with b M | b and row reduce M in this case is just the 2 given vectors in the columns of M so M is 3x2 M|b is 3x3

OpenStudy (anonymous):

 I i have this ,then wat must i do?